# Topological flow data analysis for transient flow patterns: a graph-based approach

### v1

@@ 2.2. Topological Flow Data Analysis
We provide a brief review of TFDA used in this paper. For a complete description of the theory, see Yokoyama and Sakajo (2013); Sakajo and Yokoyama (2018); Uda et al. (2019). Let us consider a Hamiltonian vector field of $C^{r}$-class ($r\geq 1$) in a two-dimensional disk $\mathcal{D}$ satisfying the slip boundary condition. Then the particle orbits generated by this Hamiltonian vector field coincide with the level curves of the Hamiltonian function on $\mathcal{D}$. Since the streamfunction of the two-dimensional incompressible vector field is the Hamiltonian function of the flow, the level curves are called streamlines. In general, since the vector field can depend on time, the particle orbits advected by the vector field are no longer equivalent to the level curves of the Hamiltonian function. However, TFDA deals with instantaneous velocity fields and their generating particle orbits, i.e., streamlines, for every fixed time.
@@
Here, we restrict our attention to a special class of the vector field, called structurally stable Hamiltonian vector fields. By structural stability, we mean that the topological structure of particle orbits is unchanged under any small perturbation in the $C^{r}$ topology. See Ma and Wang (2005); Yokoyama and Sakajo (2013) for a mathematically rigorous definition of structural stability. The set of structurally stable Hamiltonian vector fields in the disk $\mathcal{D}$ is denoted by $\mathcal{H}^{r}$. Such restriction gives rise to no serious problems theoretically since $\mathcal{H}^{r}$ is a generic subset of the set of all $C^{r}$ Hamiltonian vector fields. That is to say, any Hamiltonian vector field is approximated by a sequence of structurally stable Hamiltonian vector fields in $\mathcal{H}^{r}$. In addition, practically, numerical simulation of flows mostly generates structurally stable streamline patterns since it hardly resolves structurally unstable orbits due to numerical approximation errors.
@@
Figure 4. Local orbit structures appearing in the streamline patterns of structurally stable Hamiltonian vector fields in $\mathcal{D}$. (a) Figure-eight patterns with a saddle and two self-connected saddle separatrices whose COT symbol is $b_{\pm\pm}$. (b) Local orbit structures with a saddle, in which one saddle connection encloses another. The COT symbol is $b_{\pm\mp}$. (c) Saddle connections between two different saddles on the same boundary. The COT symbol is $c_{\pm}$. (d) Isolated boundaries, $\beta_{\pm}$. (e) Elliptic centres, $\sigma_{\pm}$. In each panel, $\Box_{b_{\pm}}$ and $\Box_{c_{\pm}}$ indicate that the local orbit structures in (6) are embedded in the flow structures as their internal structure.
@@
To describe the theory of TFDA, we introduce all local orbit structures that appear in structurally stable Hamiltonian flows in $\mathcal{D}$. We then provide a unique letter, called a COT symbol, to each local orbit structure. Figure 4(a) shows a figure-eight orbit structure with two self-connected saddle connections. To these local orbit structures, we assign the COT symbols $b_{++}$ and $b_{--}$, in which the subscript $\pm$ indicates the direction of the two self-connected saddle connections. Namely, we use the subscript $+$ for the saddle connection in the anti-clockwise direction and the subscript -- for the clockwise one. We have other local orbital structures as in Figure 4(b) consisting of a saddle and two self-connected saddle connections, in which one saddle connection surrounds another. Since the flow directions along the saddle connections become the opposite, its COT symbol is given by $b_{\pm\mp}$. Figure 4(c) shows local orbit structures consisting of a saddle connection between two different saddles on the same boundary. The COT symbol $c_{\pm}$ is assigned to the structures according to the orientations of the saddle connections. We further consider an isolated boundary along which the flow is going in the anti-clockwise (resp. clockwise) direction in Figure 4(d). The COT symbol is given by $\beta_{+}$ (resp. $\beta_{-}$). The box symbols $\Box_{b_{\pm}}$ and $\Box_{c_{\pm}}$ in Figure 4(a–d) express that local orbit structures are embedded as internal structures there. In other words, each box symbol corresponds to a local orbit structure chosen as
@@
For instance, in the local orbit structures of $c_{\pm}$ in Figure 4(c), the saddle connection needs to enclose one local orbit structure from (6) in $\Box_{b_{\pm}}$ as an internal structure. On the other hand, one can attach any number of the local orbit structures $\Box^{i}_{c_{\pm}}$ ($i=1,\dots n$) at regions enclosed by dashed curves along the boundary, but these structures are not necessarily present, i.e. $n=0$ is allowed. The same situation also applies to the structure $\beta_{\pm}$. Finally, we define an isolated elliptic centre whose COT symbol is $\sigma_{+}$ (resp. $\sigma_{-}$) associated with anti-clockwise (resp. clockwise) periodic orbits in Figure 4(e). Note that they have no internal structures.
Figure 5. Root structures in the disk $\mathcal{D}$, where a local orbit structures $b_{\pm\pm}$ or $b_{\pm\mp}$ can be embedded and any number of $c_{\pm}$ structures are attached to the boundary of the disk. (a) The root structure $\beta_{\emptyset+}$. The flow along the boundary is going in the anticlockwise direction. (b) The root structure $\beta_{\emptyset-}$. The flow along the boundary is going in the clockwise direction.
For the topological orbit structures generated by structurally stable Hamiltonian vector fields in $\mathcal{D}$, the following result has been shown by Yokoyama and Sakajo (2013).
@@
The topological structure of particle orbits (streamlines) generated by structurally stable Hamiltonian vector field in $\mathcal{D}$ is uniquely represented by a combination of the local orbit structures $b_{\pm\pm}$, $b_{\pm\mp}$, $c_{\pm}$, $\sigma_{\pm}$ and $\beta_{\pm}$.
Based on the classification theorem, we convert the topological structure of orbit connections for a given structurally stable Hamiltonian vector field in $\mathcal{D}$ into a planar tree and its string expression, which are the output of TFDA. The conversion algorithm is outlined as follows. For a detailed description of the algorithm, refer to Uda et al. (2019). Let us first consider two topologically simplest flows in the disk $\mathcal{D}$ as shown in Figure 5, which are called the root structures. When the flow direction along the boundary of the disk is anti-clockwise (resp. clockwise), we assign the COT symbol $\beta_{\emptyset+}$ (resp. $\beta_{\emptyset-}$). They indicate that the root structure contains at least one local orbit structure in $\Box_{b_{\pm}}$ and any number of local orbit structures at $\Box_{c_{\pm}}$ can be attached along the boundary. Next, for the given structurally stable Hamiltonian vector field (or the Hamiltonian function), starting from the root structures, we identify the internal local orbit structures corresponding to $\Box_{b_{\pm}}$ and $\Box_{c_{\pm}}$ inductively. Every time we identify internal local orbit structures, we create an edge from the parent node of the structure containing them to the child nodes of embedded internal local structures. Repeating this step inductively until we reach isolated elliptic centres or boundaries with no $c_{\pm}$ structures attached yields a planer tree, namely a partially Cyclically Ordered rooted labelled Tree (COT). According to Sakajo and Yokoyama (2018), it has been mathematically shown that if the root node is fixed, COTs are in one-to-one correspondence with the topological orbit structures of structurally stable Hamiltonian vector fields. This means that a COT becomes a unique identifier of the flow pattern in terms of topology. Remark that the COT is equivalent to the Reeb graph of the streamfunction (Uda et al., 2019)
@@ 2.3. TFDA for the lid-driven cavity flow
Since the flow goes around the boundary of the cavity in the clockwise direction, the root structure of this flow is identified as $\beta_{\emptyset-}$ as in Figure 5(b). As an internal structure of the cavity flow corresponding to $\Box_{b_{-}}$ of the root structure $\beta_{\emptyset-}$, we identify an elliptic centre $\sigma_{-}$ located at the centre of a large rotating flow region. For orbit structures of $\Box_{c_{+}}$ along the boundary, we find three local orbit structures $c_{+}$ in Figure 4(c) connecting between saddles on the boundary at the corners of the cavity. These orbit structures express the existence of the recirculating regions. Here, to distinguish these $c_{+}$ structures, we assign the enumerated COT symbols $c_{+}^{0}$, $c_{+}^{1}$, and $c_{+}^{2}$ to the recirculating regions the top left corner, the bottom left corner, and the bottom right corner respectively, as shown in Figure 6(a). Since the elliptic centre $\sigma_{-}$ and the three $c_{+}$ structures are identified as the internal structures contained in the root structure $\beta_{\emptyset-}$, we create the root node of $\beta_{\emptyset-}$ and four child nodes of the local orbit structures $\sigma_{-}$ and $c_{+}^{i}$ for $i=0,1,2$, and we then connect them with edges. Since the elliptic centre $\sigma_{-}$ in the centre of the rotating region has no internal structure, we stop the procedure for this node. Similarly, since only elliptic centres $\sigma_{+}$ are contained in the orbit structure $c_{+}^{0}$, the procedure stops after adding the node of the elliptic centre with the label $\sigma_{+}$ as its child node. On the other hand, since an anti-clockwise figure-eight $b_{++}$ structure in Figure 4(a) is contained inside the $c_{+}^{1}$ structure at the bottom left corner, a new node with label $b_{++}$ is created as a child node and an edge connecting to the parent node with the label $c_{+}^{1}$. Both recirculating domains $c_{+}^{1}$ and $c_{+}^{2}$ contain a small recirculating domain $c_{-}$ inside, where the flow is rotating in the opposite direction. Finally, there are isolated elliptic centres $\sigma_{+}$ inside the orbit structure $b_{++}$ and $\sigma_{-}$ inside the structure $c_{-}$, so the conversion algorithm terminates. As a result, we obtain the COT as shown in Figure 6(b) expressing the topological structure of the streamline pattern.
The COT is expressed as a sequence of COT symbols, say COT representation, which is given by
(7) $\beta_{\emptyset-}(\sigma_{-},c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(b_{++}\{\sigma_{+},\sigma_{+}\},c_{-}(\sigma_{-}))\cdot c_{+}^{2}(\sigma_{+},c_{-}(\sigma_{-})).$
It is obtained by using a regular tree grammar with cyclic order introduced in Yokoyama and Yokoyama (2020), which is a generalization of the regular tree grammar (cf. Comon et al. (2008)). The specific implementation procedure for the grammar, which is a theoretical conversion rule, is described in Uda et al. (2019). Since we use the COT symbols as the labels of the nodes of the tree, we refer to the local orbit structures with the COT symbols in the COT representations.
We perform our TFDA calculations using psiclone, a Python library that implements the algorithm computing COT representations developed by Uda et al. (2019). First, given a grid dataset representing a streamline function $\psi$, we compute the persistent homologies of sublevel and superlevel filtrations of $\psi$. The adjacency structure of the data grid is assumed to follow the von Neumann neighbourhood ($4$-neighbourhood). Given a prescribed threshold value $\varepsilon>0$, we discard intervals (or bars) in both persistent homologies whose lifetimes (or lengths) are less than $\varepsilon$. In our analysis, we set the threshold value to $\varepsilon=5.0\times 10^{-4}$, chosen to ignore small-scale variations of the streamfunction, particularly those near the boundary of the computational domain. Second, we combine both persistent homologies and obtain the Reeb graph of $\psi$, which provides a global topological summary of the nesting information of the streamlines. Third, we scan all the edges in the Reeb graph and convert their local patterns into nested COT symbols inductively, as described in the previous subsection. Here, since there are multiple choices for the COT root nodes, we fix the choice by selecting the representative root point minimising the streamline function value. That is to say, the scanning of level sets starts from this root point $\operatorname{argmin}\psi(x,y)$. After obtaining the COT, we incorporated supplementary information by assigning corner location numbers $i=0,1,2$ to $c_{+}$ in the COT, representing the recirculating flow domains at the top left, bottom left, and bottom right corners, respectively. The enumeration of the COT symbol $c_{+}$ for the top right corner is omitted, as the lid-driven cavity configuration does not allow for large-scale local orbit structures in that region. Note that the number of child nodes of type $c_{+}$ varies over time, and a corner may have multiple $c_{+}$ nodes or none at all.
3. Numerical Results
3.1. Detailed analysis at $Re=14000$
3.1.1. Transition of streamline topology

### v2

@@ 2.2. Topological Flow Data Analysis
We provide a brief review of TFDA used in this paper. For a complete description of the theory, see Yokoyama and Sakajo (2013); Sakajo and Yokoyama (2018); Uda et al. (2019). We consider a Hamiltonian vector field of $C^{r}$-class ($r\geq 1$) in a two-dimensional disk $\mathcal{D}$ satisfying the slip boundary condition. Then the particle orbits generated by this Hamiltonian vector field coincide with the level curves of the Hamiltonian on $\mathcal{D}$. Since the streamfunction of the two-dimensional incompressible vector field is the Hamiltonian of the flow, the level curves are called streamlines. As explained in the introduction, since the vector field can depend on time, the particle orbits advected by the vector field are no longer equivalent to the level curves of the Hamiltonian in general. However, TFDA deals with velocity fields and their generating particle orbits for every fixed time, i.e., instantaneous streamlines.
@@
Here, we restrict our attention to a special class of vector fields, called structurally stable Hamiltonian vector fields. By structural stability, we mean that the topological structure of instantaneous streamlines is unchanged under any small perturbation in the $C^{r}$ topology. See Ma and Wang (2005); Yokoyama and Sakajo (2013) for a mathematically rigorous definition of structural stability. The set of structurally stable Hamiltonian vector fields on the disk $\mathcal{D}$ is denoted by $\mathcal{H}^{r}$. This restriction gives rise to theoretically no serious problems, since $\mathcal{H}^{r}$ is a generic subset of the set of all $C^{r}$ Hamiltonian vector fields. That is, any Hamiltonian vector field is approximated by a sequence of structurally stable Hamiltonian vector fields in $\mathcal{H}^{r}$. In addition, practically, numerical simulation of flows mostly generates structurally stable streamline patterns since it hardly resolves structurally unstable orbits due to numerical approximation errors.
@@
Figure 4. Local orbit structures appearing in the streamline patterns of structurally stable Hamiltonian vector fields in $\mathcal{D}$ and their COT symbols. (a) Figure-eight patterns with a saddle and two self-connected saddle separatrices whose COT symbol is $b_{\pm\pm}\{\Box_{b_{\pm}},\Box_{b_{\pm}}\}$. (b) Local orbit structures with a saddle, in which one saddle connection encloses another. The COT symbol is $b_{\pm\mp}(\Box_{b_{\pm}},\Box_{b_{\mp}})$. (c) Saddle connections between two different saddles on the same boundary. The COT symbol is $c_{\pm}(\Box_{b_{\pm}},\Box_{c_{\mp}}^{1},\ldots,\Box_{c_{\mp}}^{n})$. (d) Isolated boundaries $\beta_{\pm}\{\Box_{c_{\pm}}^{1}\cdot\cdots\cdot\Box_{c_{\pm}}^{n}\}$. (e) Elliptic centres $\sigma_{\pm}$. In each panel, $\Box_{b_{\pm}}$ and $\Box_{c_{\pm}}$ indicate that the local orbit structures in (6) are embedded in the flow structures as their internal structure. The curly brackets $\{\}$ and the round brackets $()$ in COT symbols express whether the arrangement of the embedded structures is determined up to cyclic order and is uniquely determined respectively.
@@
To describe the theory of TFDA, we introduce all local orbit structures that appear in structurally stable Hamiltonian flows in $\mathcal{D}$. We then provide a unique letter, called a COT symbol, to each local orbit structure. Figure 4(a) shows a figure-eight orbit structure with two self-connected saddle connections. For these local orbit structures, we assign the COT symbols $b_{++}\{\Box_{b_{+}},\Box_{b_{+}}\}$ and $b_{--}\{\Box_{b_{-}},\Box_{b_{-}}\}$. The subscript $\pm$ indicates the direction of the two self-connected saddle connections. That is, we use the subscript ++ for the saddle connection in the anticlockwise direction and the subscript - for the clockwise one. The box symbols $\Box_{b_{\pm}}$ represent the local orbit structures enclosed by the saddle connections, and the curly brackets $\{\}$ indicate that the arrangement of the inner local orbit structures is determined to be in cyclic order. The local orbit structures in $\Box_{b_{\pm}}$ can be chosen as follows.
@@
We have other local orbital structures as in Figure 4(b) consisting of a saddle and two self-connected saddle connections, in which one saddle connection surrounds another. Since the flow directions along the saddle connections become the opposite, its COT symbol is given by $b_{\pm\mp}(\Box_{b_{\pm}},\Box_{b_{\mp}})$, in which the round brackets $()$ mean that the arrangement of the inner local orbit structures is determined uniquely. Figure 4(c) shows local orbit structures consisting of a saddle connection between two different saddles on the same boundary. The COT symbol $c_{\pm}(\Box_{b_{\pm}},\Box_{c_{\mp}}^{1},\ldots,\Box_{c_{\mp}}^{n})$ is assigned to the local orbit structures according to the orientations of the saddle connections, indicating that a finite number of local orbit structures $\Box_{c_{\pm}}^{s}$, $s=1,\dots,n$, chosen as (6), can be attached to the boundary in the anticlockwise direction if there exist. We also consider an isolated boundary along which the flow is going in the anticlockwise (resp. clockwise) direction in Figure 4(d). The COT symbol is given by $\beta_{+}\{\Box_{c_{+}}^{1}\cdot\,\cdots\,\cdot\Box_{c_{+}}^{n}\}$ (resp. $\beta_{-}\{\Box_{c_{-}}^{1}\cdot\,\cdots\,\cdot\Box_{c_{-}}^{n}\}$), in which a finite number of local orbit structures are attached in the anticlockwise direction if exists, and they are arranged with the dot $\cdot$” within $\{\}$ to designate the multiple cyclic order arrangement. Finally, we define an isolated elliptic centre whose COT symbol is $\sigma_{+}$ (resp. $\sigma_{-}$) associated with anticlockwise (resp. clockwise) periodic orbits in Figure 4(e). Note that they have no internal structures.
Figure 5. (a)An anticlockwise recirculating region on the boundary that is represented by $c_{+}(\sigma_{+})$. (b) An isolated anticlockwise recirculating region represented by $b_{+-}(\sigma_{+},\sigma_{-})$.
Figure 6. Root structures in the disk $\mathcal{D}$ and their COT symbols, where a local orbit structure $_{\pm\pm}, b±b_{\pm\mp}$, $\sigma_{\pm}$ or $\beta_{\pm}$ is embedded at $\Box_{b_{\pm}}$ and any number of $c_{\pm}$ structures can be attached along the boundary of the disk as in $\Box_{c_{\pm}}^{s}$, $s=1,\dots,n$ in the anticlockwise cyclic order if exist. (a) The root structure $\beta_{\emptyset+}$. The flow along the boundary is going anticlockwise direction. (b) The root structure $\beta_{\emptyset-}$. The flow along the boundary is going clockwise direction.
For the topological orbit structures generated by structurally stable Hamiltonian vector fields in $\mathcal{D}$, the following result is shown by Yokoyama and Sakajo (2013).
@@
The topological structure of instantaneous streamlines generated by a structurally stable Hamiltonian vector field in $\mathcal{D}$ is uniquely represented by a combination of local orbit structures $_{\pm\pm}, b±b_{\pm\mp}$, $c_{\pm}$, $\sigma_{\pm}$ and $\beta_{\pm}$.
We note that a change of the COT symbols can provide specific hydrodynamic information. For example, an anticlockwise boundary saddle represented by $c_{+}(\sigma_{+})$ in Figure 5(a) can transition to a pattern with a homoclinic saddle expressed as $b_{+-}(\sigma_{+},\sigma_{-})$ in Figure 5(b). This transition represents a physical mechanism where an anticlockwise vortical flow region, confined to the boundary, separates from the boundary and enters the interior of the fluid region, i.e., a vortex separation. More examples are demonstrated in Sakajo et al. (2014); Sakajo and Yokoyama (2015).
Based on the classification theorem, we convert the topological structure of orbit connections for a given structurally stable Hamiltonian vector field in $\mathcal{D}$ into a planar tree (COT) and its string expression (COT representation), which are the output of TFDA. The conversion algorithm to COT is described below. For a detailed description of the algorithm, see Appendix B and Uda et al. (2019). Let us first consider two topologically simplest flows in the disk $\mathcal{D}$ as shown in Figure 6, which are called root structures. When the flow direction along the boundary of the disk is anticlockwise (resp. clockwise), we assign the COT symbol $\beta_{\emptyset+}(\Box_{b_{+}},\{\Box_{c_{-}}^{1}\cdot\,\cdots\,\cdot\Box_{c_{-}}^{n}\})$ (resp. $\beta_{\emptyset-}(\Box_{b_{-}},\{\Box_{c_{+}}^{1}\cdot\,\cdots\,\cdot\Box_{c_{+}}^{n}\})$). They indicate that the root structure contains at least one local orbit structure in $\Box_{b_{\pm}}$ and any number of local orbit structures $\Box_{c_{\pm}}^{s}$, $s=1,\dots,n$ can be attached along the boundary in anticlockwise cyclic order if there exist. Next, for the given structurally stable Hamiltonian vector field (or the Hamiltonian), starting from the root structure, we identify the internal local orbit structures corresponding to $\Box_{b_{\pm}}$ and $\Box_{c_{\pm}}^{s}$ inductively. Every time we identify internal local orbit structures, we create edges from the parent node of the structure containing them to the child nodes of embedded internal local structures. Repeating this step inductively until we reach isolated elliptic centres or boundaries with no $c_{\pm}$ structures attached yields a planar tree, namely a partially Cyclically Ordered rooted labelled Tree (COT). According to Sakajo and Yokoyama (2018), it has been mathematically shown that if the root node is fixed, COTs are in one-to-one correspondence with the topological orbit structures of structurally stable Hamiltonian vector fields. In other words, each topologically different pattern of instantaneous streamlines maps to exactly one COT, and vice versa up to topological equivalence. This means that a COT becomes a unique identifier of the flow pattern in terms of topology.
@@ 2.3. TFDA for the lid-driven cavity flow
We show that the topological structure of the streamline pattern in Figure 7(a) is converted to the COT in Figure 7(b). First, since the flow goes around the cavity boundary in the clockwise direction, the root structure of this flow is topologically identified as $\beta_{\emptyset-}$ in Figure 6(b). As an internal structure of the root structure corresponding to $\Box_{b_{-}}$, we identify an elliptic centre $\sigma_{-}$ located at the centre of a large rotating flow region. For orbit structures of $\Box_{c_{+}}^{s}$ along the boundary, we find three local orbit structures $c_{+}$ connecting between saddles at the boundary at the corners of the cavity. These orbit structures express the existence of the recirculating regions. Here, to distinguish these $c_{+}$ structures, we tag each of them with a corner index $i$ like $c_{+}^{i}$ that indicates which corner of the cavity it belongs to (see Figure 7(a)):
(7) $i=\begin{cases}0&\text{(the top left)},\\ 1&\text{(the bottom left)},\\ 2&\text{(the bottom right)}.\end{cases}$
The top right corner index is excluded because the moving lid suppresses large persistent orbits there, as explained later. At any given time, a corner may host zero, one, or several $c_{+}$ symbols, reflecting the intermittent birth and death of recirculation cells. Since the elliptic centre $\sigma_{-}$ and the three $c_{+}$ structures are identified as the internal structures contained in the root structure $\beta_{\emptyset-}$, we create the root node of $\beta_{\emptyset-}$ and four child nodes of the local orbit structures $\sigma_{-}$ and $c_{+}^{i}$ for $i=0,1,2$, and then connect them with the edges. Second, since the elliptic centre $\sigma_{-}$ in the centre of the rotating region does not have an internal structure, we stop the procedure for this node. Similarly, since only elliptic centres $\sigma_{+}$ are contained in the orbit structure $c_{+}^{0}$, the procedure is stopped after adding an edge to the node of the elliptic centre with the label $\sigma_{+}$ as its child node. On the other hand, since an anticlockwise $b_{++}$ structure in Figure 4(a) is contained inside the $c_{+}^{1}$ structure in the bottom left corner, a new node with the label $b_{++}$ is created as a child node and an edge connecting to the parent node with the label $c_{+}^{1}$. Both recirculating domains $c_{+}^{1}$ and $c_{+}^{2}$ contain a small recirculating domain $c_{-}$ inside, where the flow rotates in the opposite direction. Finally, there are isolated elliptic centres $\sigma_{+}$ within the orbit structure $b_{++}$ and $\sigma_{-}$ within the structure $c_{-}$, so the conversion algorithm is terminated.
The COT is compactly described as the following string, say COT representation.
(8) $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(b_{++}\{\sigma_{+},\sigma_{+}\},c_{-}(\sigma_{-}))\cdot c_{+}^{2}(\sigma_{+},c_{-}(\sigma_{-}))\}).$
The conversion from the COT of Figure 7(b) to (8) is explained below. We begin with the root node $\beta_{\emptyset-}$ whose COT symbol is represented by $\beta_{\emptyset-}(\Box_{b_{-}},\{\Box_{c_{+}}^{1}\cdot\,\cdots\,\cdot\Box_{c_{+}}^{n}\})$. Since the root node has four child nodes, by substituting the COT symbols for the nodes into the positions of $\Box_{b_{-}}$ and $\Box_{c_{+}}^{s}$ for $s=1,2,3$ as $\Box_{b_{-}}=\sigma_{-}$, $\Box_{c_{+}}^{1}=c_{+}^{0}(\Box_{b_{+}},\Box_{c_{-}}^{1},\ldots,\Box_{c_{-}}^{n})$, $\Box_{c_{+}}^{2}=c_{+}^{1}(\Box_{b_{+}},\Box_{c_{-}}^{1},\ldots,\Box_{c_{-}}^{n})$ and $\Box_{c_{+}}^{3}=c_{+}^{2}(\Box_{b_{+}},\Box_{c_{-}}^{1},\ldots,\Box_{c_{-}}^{n})$, we obtain
$\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\Box_{b_{+}},\Box_{c_{-}}^{1},\ldots,\Box_{c_{-}}^{n})\cdot c_{+}^{1}(\Box_{b_{+}},\Box_{c_{-}}^{1},\ldots,\Box_{c_{-}}^{n})\cdot c_{+}^{2}(\Box_{b_{+}},\Box_{c_{-}}^{1},\ldots,\Box_{c_{-}}^{n})\}).$
Since the node $c_{+}^{1}$ has two nodes $b_{++}$ and $c_{-}$, we substitute $\Box_{b_{+}}=b_{++}\{\Box_{b_{+}},\Box_{b_{+}}\}$ and $\Box_{c_{-}}^{1}=c_{-}(\Box_{b_{-}})$ in the COT symbol $c_{+}^{1}(\Box_{b_{+}},\Box_{c_{-}}^{1},\ldots,\Box_{c_{-}}^{n})$ with $n=1$. Similarly, for the other nodes $c_{+}^{i}$ ($i=0,2)$, we substitute the COT symbols in $\Box_{b_{\pm}}$ and $\Box_{c_{\pm}}$ into the round brackets that follow them according to their respective order.
$\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(b_{++}\{\Box_{b_{+}},\Box_{b_{+}}\},c_{-}(\Box_{b_{-}}))\cdot c_{+}^{2}(\sigma_{+},c_{-}(\Box_{b_{-}}))\}).$
Finally, since $\sigma_{\pm}$ appear below each of $b_{++}$ and $c_{-}$ in the COT, substituting $\sigma_{\pm}$ in the corresponding to the box symbols $\Box_{b_{\pm}}$, we obtain (8) as desired.
In this paper, the computation of COT and COT representations is carried out with psiclone, a Python library that implements the TFDA algorithm proposed in Uda et al. (2019). A two-dimensional array of streamfunction values, $\psi(x,y)$, on a rectangular grid is provided as input. Psiclone extracts critical values of $\psi(x,y)$ that correspond to level sets containing saddles or centres. It is worth noting that this extraction does not require the direct calculation of separatrices, saddles, and centres. Instead, it leverages the fact that the topology of connected components enclosed by the level set changes just below and above these critical values. Indeed, persistent homolog, a technique from Topological Data Analysis (TDA), can discover all such critical values. Furthermore, this method incorporates the streamfunction values in topological change events, enabling the identification of noise-like structures if the difference in these critical values is small. In the present analysis, we keep only those topological structures that remain unchanged under perturbations of the streamfunction with amplitude $\Delta\psi<\varepsilon$ ($\varepsilon=5.0\times 10^{-4}$), thus discarding topological structures caused by small changes. In spirit, this plays the role of a low-pass filter, but the mechanism respects the robustness to topological structures, rather than frequency filtering. See Appendix B for a technical review of the algorithm.
3. Topological flow data analysis for $Re=14000$
3.1. Transition of streamline topology